OpenStudy (anonymous):

Let u = <-5, 1>, v = <7, -4>. Find 9u - 6v

OpenStudy (jdoe0001):

\(\bf u=<a,b>\qquad v=<c,d>\quad \begin{cases} 9u\to 9<a,b>\to <9\cdot a,9\cdot b>\\ \bf 6v\to 6<c,d>\to <6\cdot c,6\cdot d> \end{cases} \\ \quad \\ u-v\implies <a,b>-<c,d>\implies <a-c,b-d>\)

OpenStudy (jdoe0001):

so you'd do the "scalar multiplication" first, and whatever vectors you get, SUM them up, or SUBTRACT them in this case

OpenStudy (dan815):

now replace a and b, c and d, with the number they gave you, and follow what john doe told you...

OpenStudy (anonymous):

is it <-87, 33> ?

OpenStudy (dan815):

i only checked your x part of the vector it is right..

OpenStudy (dan815):

ok yes it is right

OpenStudy (anonymous):

Thanks!